OpenStudy (anonymous):

Hi, I have a difficult calculus type question, ive been working on it for ages, Need Help

OpenStudy (anonymous):

Attached is the question

OpenStudy (dumbcow):

someone else had this very same question earlier, let me see if i can find the post

OpenStudy (anonymous):

\[ L(x) = \sqrt{(0.2-x)^2+(2.7-y)^2} = \sqrt{(0.2-x)^2+(2.7-x^2)^2} \]

OpenStudy (anonymous):

if it was me, i would get rid of that nasty square root, and work with the square of the distance instead it will make finding the derivative and differentiation tons easier

OpenStudy (dumbcow):

sorry i cant find it

OpenStudy (dumbcow):

anyway once you differentiate and set numerator equal to 0 here is how you do Newtons Method: \[\large x_{n+1} = x_n-\frac{f(x_n)}{f'(x_n)}\] f(x) is numerator of L'(x) x_0 is initial guess ... i would try x=1 you keep repeating the process until x_n+1 = x_n or f(x) = 0 it usually takes about 5 iterations...plugging formula into Excell will make it faster

OpenStudy (dumbcow):

you should end up with \[f(x) = 4x^{3}-8.8x-0.4 = 0\] \[\rightarrow f(x) =x^{3}-2.2x-0.1 = 0\] \[f'(x) = 3x^{2}-2.2\]

OpenStudy (anonymous):

@dumbcow I got the f(x) and f'(x), im trying to find out the newton's method

OpenStudy (anonymous):

And then when i have to find the global minimum, do i just use the equation equal it to zero and second derivative? or is it also through newtons?

OpenStudy (dumbcow):

well i posted Newtons method...which will give you the 3 zeros **Note** the zero you obtain depends on the initial guess, make sure they are far apart as far as global min, its just the smallest value of L(x) after you plug in the 3 zeros obtained from Newtons method

OpenStudy (anonymous):

Oh thanks

OpenStudy (dumbcow):

also the diagram gives you a hint, obviously the point is somewhere around x=1.5 so the global min is value from zero closest to 1.5

OpenStudy (anonymous):

I posted this solution yesterday

OpenStudy (dumbcow):

@CarlosGP , very nicely done thanks for posting it again

OpenStudy (anonymous):

My pleasure!

OpenStudy (anonymous):

Here is yesterday's link. The thread can be useful for further clarification http://openstudy.com/users/carlosgp#/updates/523f95dfe4b0900154177a8c